OpenStudy (anonymous):

(5.1 x10^7)(6.1 x10^-2) Multiply Write using scientific notation

OpenStudy (anonymous):

Multiply 5.1 by 6.1 then multiply this by 10^{7-2} = 10^5 then make this into scientific notation once more, 31.11 times 10^5 = 3.111 times 10^6

Directrix (directrix):

(5.1 x10^7)(6.1 x10^-2)= (5.1)(6.1) * 10 ^ (7 -2) = 31.11 * 10 ^ 5 = 3.111 * 10^6